OpenStudy (anonymous):

Rationalize the denominator. The answer should be expressed in simplified form.

OpenStudy (anonymous):

\[\frac{ \sqrt{10}-\sqrt{7} }{ \sqrt{10}+\sqrt{3} }\] @charliealbert

OpenStudy (anonymous):

\[\frac{ \sqrt{10}-\sqrt{3} }{ \sqrt{10}-\sqrt{3} }*\frac{ \sqrt{10}-\sqrt{7} }{ \sqrt{10}+\sqrt{3} }\] \[\frac{ 10-\sqrt{10}\sqrt{7}-\sqrt{3}\sqrt{10}+\sqrt{7}\sqrt{3} }{ 10-3 }\]

OpenStudy (anonymous):

In the numerator: (a+b)(c+d)=a(c+d)+b(c+d)

OpenStudy (anonymous):

Do the 10s divide out?

OpenStudy (anonymous):

@charliealbert

OpenStudy (anonymous):

What do you mean?.

OpenStudy (anonymous):

Does the 10 in the numerator and denominator cancel out?

OpenStudy (anonymous):

No because 3 is subtracting 10.

OpenStudy (anonymous):

the anwer would be: \[\frac{ 10-\sqrt{70}-\sqrt{30}+\sqrt{21} }{ 7 }\]

OpenStudy (anonymous):

Oh ok thank you! Sorry i took so long I was registering for my classes for the fall online. :)

OpenStudy (anonymous):

Ok, no problem. Good luck with your autumn classes

OpenStudy (anonymous):

Thanks!